PM.62 The US government has recently contracted with Boeing to build its fleet of airborne tankers. The project has been laid out with durations (in weeks), predecessors, and costs included:

Activity   Normal Time (weeks)   Crashed Time (weeks)   Normal Cost   Crashed Cost   Predecessor(s)  
A 11 10 $1,500 $1,650 -
B 7 5 $1,200 $2,490 -
C 5 4 $900 $1,200 B
D 4 2 $2,100 $3,000 A
E 2 1 $550 $1,250 B
F 4 3 $660 $1,050 E
G 10 6 $560 $1,040 C, D, F

What is the critical path of the project?

B-C-G

A-D-G

B-E-F-G

If you wanted to shorten the overall duration of the project, which activity would you crash first?

B

E

C

G

What would be the weekly crashing cost of the activity you selected in the previous question? (Display your answer to two decimal places.)

   

Crash the project by 5 weeks. What is the total cost of crashing the project (just the additional cost, not total project cost)? (Display your answer to two decimal places.)  
Answer #1

Critical path is longest path of the project = A-D-G = 25 weeks

We will crash activity G first as it is on critical path.

Weekly crashing cost = (1040-560) /(10-6) = $120 / week

Now to crash project by 5 weeks, we first crash G by 4 weeks. This shortens all paths by 4 weeks. Next we crash A by 1 week. The project duration now is 20 weeks. We Crash A as it has lower weekly crash cost compared to D.

Total additional cost = Crash cost - Normal cost = (1040-560) +(1650-1500) = $630